OpenStudy (anonymous):

A hot tub holding 480 gal of water begins to empty at a rate of 5 gal/min. At the same time, an empty hot tub begins to fill at a rate of 15 gal/min. Let g represent the number of gallons of water and let t represent time in minutes. The system models this situation. How long will it take for the hot tubs to hold equal amounts of water, and how much water will that be? g = 480 – 5t g = 15t It will take____minutes for both hot tubs to hold equal amounts of water. They will each hold_____ gallons.

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Plug in 15t for g. 15t=480-5t 20t=480 t=24 g=15(24) g=360